# CAPTURe: Evaluating Spatial Reasoning in Vision Language Models via Occluded Object Counting

### v1

@@ 4.2 Effect of Data Factors on VLM Performance
Here, we use the CAPTURe data (which can be controlled precisely and for which we have metadata) to examine which features correlate with model performance. We test the effect of the following variables on final performance: (1) Increasing the number of occluded objects; (2) Varying the pattern. We also investigate whether models are biased towards certain numbers, whether models can classify patterns, and to what degree models can predict the number of occluded objects only (rather than the total).
Models perform worse when more dots are occluded.
In Fig. 6 (right), we observe that error increases with respect to the number of occluded dots. However, Fig. 6 (left) also shows that performance is less affected by the total number of dots. This suggests that the task difficulty is more closely correlated with the difficulty of occlusion – i.e. the difficulty of the world modeling task – rather than the complexity of the pattern. However, there are models that do not follow this trend, e.g. GPT-4o, which has lower error on certain numbers. We investigate this further below.
Models are more likely to predict certain numbers.
To examine where models frequently err, we generated a confusion matrix for every model based on CAPTURe results (shown in Sec. C.4). We find that models often over-predict numbers associated with common counts in real life: GPT-4o tends to predict numbers like 8, 9, 10, and 12, which are all non-prime numbers (i.e. can be arranged into a grid) and common groupings of objects, e.g. 12 is a common grouping (dozens) and allows arrangements into 3x4 or 2x6 grids. InternVL and Qwen2VL over-predict 5 and 10, aligning with how humans conceptualize numbers. Indeed, Coupland [12] found that numbers 5, 10, 20, and other round numbers appear disproportionally more in online texts. Molmo shows no correlation with these factors, possibly because of its unique “point and count” ability.
Performance depends on pattern type.
The controllability of CAPTURe allows us to measure the effect of pattern type on performance. In Fig. 7, we find that model performance differs across shapes with some regularity: objects arranged in a circle generally have lower sMAPE than other shapes, across all models.
@@
Models can identify patterns.
To determine how much model errors can be attributed to a lack of pattern recognition ability, we formulate a separate task where models must recognize the pattern in the image on CAPTURe. Here, we frame the task as a multiple-choice task, asking the model to select from the pattern types available (rectangle, triangle, or circle). Table 4 illustrates that all perform substantially better than random at this task, with most models except InternVL2 achieving accuracy above 80% in the unoccluded setting. As expected, the patterns were easier to identify in unoccluded scenarios, with models suffering a 10.95% accuracy drop on average in the occluded setting. Notably, GPT-4o and Qwen2VL have a fairly small drop in performance, suggesting they are generally able to capture the pattern even in the presence of occlusion.
@@ 4.3 Analysis with Auxiliary Information
In Sec. 4.1, we see that models broadly struggle with amodal counting. Here, we seek to disentangle whether this problem results from a failure to reason, the absence of a world model, or both, by giving VLMs two different types of auxiliary information: oracle information and predicted information. Oracle information is ground truth and is directly pulled from CAPTURe’s metadata, e.g., object locations. Predicted information generates new information from a completely separate model and gives it to the VLM. This information is not ground truth and is sourced from an external model, such as an image inpainting model, rather than the VLM. By giving the model auxiliary information in the form of reasoning and spatial clues, we can establish how much of each model’s error results from an inability to handle occlusion rather than an inability to recognize and count visible objects.
Oracle setup.
@@
We test two oracles for CAPTURe’s occluded split based on its constituent subtasks: counting the visible objects and inferring/counting occluded objects. Both oracles provide the VLM with text-based coordinates of objects in the image, simplifying the visual task by assuming the VLM effectively has a perfect visual system that can recognize and localize objects in the image. The first oracle, the Visible Object Coordinate Oracle, gives VLM the coordinates of all unoccluded objects (encoded as text, as seen in Fig. 8) and instructs the model to estimate the number of occluded objects, then count the number of visible object coordinates, and add the two. In other words, the model is given oracle information about what objects are visible, thus also revealing key information about the pattern. The second oracle, the All Object Coordinate Oracle, instead gives the model the coordinates of all objects. Here, the model only needs to count the coordinates in the prompt, eliminating the need to reason on the visual input.222Note that Molmo is excluded in these tests because it contains a prompt limit that would truncate the list of coordinates. An example of the oracle inputs can be seen in Fig. 8.
Prediction setup.
@@
In this setting, we provide the VLM with an external world model representation predicted by another model. Specifically, we develop the Inpainting Pipeline to fill in the occluded region via a diffusion-based inpainting model and pass the inpainted image to the VLMs. For the inpainting model, we choose FLUX.1-Fill [dev], whose backbone FLUX.1 [dev] [21] is a top public model in the Text to Image Model Arena [7]. An example input to VLM can be seen on the far-right of Fig. 8.
Providing visible or all object coordinates improves performance substantially.
@@
The results in Tab. 2 indicate that models struggle on CAPTURe, which requires identifying a pattern and counting both visible and occluded objects. Moreover, models generally struggle with counting even in unoccluded settings. Both oracles simplify the counting task: All Object Coordinate Oracle reduces the task to simply counting coordinates with no reasoning involved, and Visible Object Coordinate Oracle similarly simplifies the task for visible objects, while still requiring inferring occluded objects. Additionally, under Visible Object Coordinate Oracle, recognizing the pattern shifts from a visual reasoning task to an augmented math problem. Instead of visually reasoning about where objects are located, the VLM considers what patterns the coordinates could make. Translating this task into a text problem results in an average increase of 15% with all objects coordinate oracle; the errors LLMs make here are due to an inability to count in the text prompt, as opposed to weaknesses in handling occlusion (since all object coordinates are given), and the strongest model, GPT-4o, achieves minimal error here. We also obtain an average increase of 8% with the visible objects coordinate oracle (shown in Tab. 5), possibly because it allows the more powerful LLM backbone (which is far larger than the vision model in all models tested) to take over and complete the counting portion of the task. Taken together, these results suggest that there is much room for improvement in visual world modeling beyond text-based reasoning of VLMs.
Providing diffusion-based inpainting improves performance marginally.
@@
Similar to the object coordinate oracles, the Inpainting Pipeline (rightmost columns in Fig. 8 and Tab. 5) eliminates the need for world modeling and provides VLMs with an approximation of the image behind the occluder. With the inpainted images, VLM error decreases by almost 2% for InternVL2 and 7% for Qwen2VL compared to the original occluded images. GPT-4o’s error increases on inpainted images by a small margin; we hypothesize that this may be because GPT-4o has one of the better world models (based on its superior performance), and thus does not improve further with the inpainted images. Moreover, every VLM still falls short of unoccluded image performance, indicating that the diffusion model is not a perfect world model. Qualitatively, we find that the inpainting model does sometimes fail to output the correct pattern behind the occluder.
@@ 5 Related Work
Spatial reasoning in visual question answering.
Past work measures the spatial reasoning capabilities of VLMs in the form of visual question answering (VQA) [4, 16] benchmarks. SpartQA [26] asks VLMs to identify the spatial relation (e.g., above, behind, left of) between objects in synthetically created 2D images from NLVR [39]. More recent benchmarks test similar spatial relation understanding with real images [24, 36, 2]. While this past work asks models to provide a text description for a relation between two fully-observed objects, CAPTURe measures the world modeling from a partially observed scene, thus requiring the handling of occlusion, pattern recognition, and counting.
Amodal completion.
Occlusions are common in natural scenes, and vision solutions for amodal completion have made significant progress infilling the occluded portions of objects [45, 38, 6]. The amodal completion task has evolved from simply completing a shape to filling in appearance (e.g., texture, color, etc.) to finally dealing with fine-grained order perception (where the model is given multiple stacked layers of occluded objects) [5]. Specifically in Qiu and Di [34], VLMs classify the hidden objects and extract fine details from occluded items. CAPTURe, however, presents a unique category of patterned amodal counting which requires inferring completely occluded objects based on a pattern rather than inferring occluded object wholes based on object parts. In other words, previous work by has only attempted tasks that require amodal completion for one object at a time [45, 31, 38], whereas CAPTURe involves inferring information about multiple objects. Multi-object amodal completion is crucial because in visual environments with lots of objects, it is likely that entire groups of objects can be occluded. Moreover, the output space of CAPTURe is language (rather than filling pixels).
Counting with vision-and-language models.
Within the task of counting, the most similar application to CAPTURe is dense counting, where the objects to be counted occlude each other. There are many practical applications of such a task, like counting microscopic cells on a crowded slide [8], determining crop yields from densely-packed fields [42], or crowd counting [43, 46, 14]. Liang et al. [23] attempted to improve crowd counting with an augmented CLIP [35], i.e. also using VLMs for counting. Additionally, Jenkins et al. [18] created a benchmark that does involve amodal counting, presenting an occluded 3D counting task where models must count objects on retail shelves. However, our work differs in many ways, as Jenkins et al. [18] only count retail shelves and processes LiDAR scans as an additional input. More broadly, dense counting focuses on overlapping objects rather than on counting objects that are arranged into patterns, which is the focus of CAPTURe.
@@ 6 Conclusion
We introduced CAPTURe, a novel benchmark for amodal counting that measures spatial reasoning capabilities under occlusion. CAPTURe is designed to assess VLMs’ ability to form a robust world model and use that model for visual reasoning skills under occlusion. By testing counting, we cast the problem as a measurable task with an objective correct answer that also has real-world utility as VLMs become more broadly adopted. Our results suggest that VLMs struggle to combine reasoning, counting, and world modeling with low performance on occluded and unoccluded images. Our analysis indicates that models improve substantially with oracle information about visible objects (simplifying the reasoning/counting tasks) and predicted information about the occluded objects (also simplifying world modeling), pointing to possible directions of model improvement.
@@ A.1 Metric Details
We use symmetric mean percent error (sMAPE) as the primary metric for our benchmarks due to its resistance to bias for under/over predictions and small/large ground truths [25]. The standard metric to use for a counting benchmark is mean average error (MAE). MAE is popular, but heavily penalizes predictions that deviate by a small margin from big ground truths, highlighting the necessity for a metric that gives equal weighting to all questions. Mean average percent error (MAPE) initially seems appealing but is disproportionally inflated for small ground truths and is biased towards overpredictions. Mean square error (MSE) and root mean square error (RMSE) are also commonly used but are very sensitive to outliers because they square the error. Intuitively, performing well on almost all questions and poorly on a small subset should score better than consistently being wrong. Among commonly-used metrics, sMAPE is the only metric that evaluates performance in relation to the distribution of ground truth elements [11]. There are two common definitions [15] for sMAPE, but we use the one that scales to 100%. sMAPE is given by:
@@ A.2 Output Tokens
To maximize the model’s chance at success, we allocate a high number of output tokens such that it can generate a rationale and output. This varies per model. We give 4000 tokens to InternVL2, 2000 tokens to Molmo, and 8192 tokens to Qwen2VL. For GPT-4o, we use the default of 4096 tokens.
@@ Appendix B CAPTURe Dataset Creation Details
While FSC-147 is a strong starting point, it cannot immediately be adapted to our task. In order to make the task of amodal counting solvable, our dataset requires images with patterns in them, such that a person (or model) can infer how the pattern would continue and thus predict the total number accurately. For questions to be solvable, the dataset’s images must be filtered down to represent patterns a model or person could recognize.
Our filtering process follows two stages. First, we prompt GPT-4o to determine whether the objects were arranged in a pattern. If the model responded with “no”, the images were immediately discarded. If the model output was “yes”, the log probability of the token is stored. Empirically, we found that higher log probability values (i.e. higher confidence scores) corresponded to more well-defined patterns in the image. Thus, we use the log probabilities for filtering.
Specifically, let be the log probability of the “yes” token and denote the threshold for determining how well-defined a pattern is.333We set based on manual evaluation, finding it resulted in fewer false positives. To filter the images based on pattern rigidity, we apply the following condition: . This inequality yields 991 images from the original dataset (16.12%). Next, we manually filter each of the selected images to ensure that the they indeed contain patterns and feature a countable number of objects, excluding 34 images. Afterward, we manually place a “fair” occluding box in each image, i.e. a box that leaves sufficient portions of the pattern visible, such that the pattern can still be inferred from the unoccluded portions of the image. Occluding boxes were also chosen with varying positions and sizes in the image.
@@ Appendix C Additional Analysis
C.1 Temperature Backoff
Model | sMAPE () | |||||||
@@ C.1 Temperature Backoff
Occluded | Unoccluded | Occluded | Unoccluded | |||||
@@
GPT-4o | 14.75 | 14.39 () | 13.34 | 12.57 () | 9.71 | 9.23 () | 5.90 | 5.93 () |
InternVL2 | 32.90 | 32.37 () | 26.17 | 27.09 () | 17.57 | 16.24 () | 16.44 | 15.59 () |
Molmo | 32.49 | 28.17 () | 25.90 | 21.23 () | 17.73 | 15.85 () | 8.40 | 2.88 () |
Qwen2VL | 29.33 | 28.47 () | 18.96 | 19.40 () | 11.74 | 11.51 () | 6.63 | 6.66 () |
Avg. of 4 VLMs | 27.37 | 25.85 () | 21.09 | 20.07 () | 14.19 | 13.21 () | 9.34 | 7.76 () |
@@
To improve VLM performance on CAPTURe, we address a trend that we established during early testing. Most of the time, the VLM fails by reaching an incorrect answer. Sometimes, however, our benchmark can cause VLMs to produce a long and irrelevant response that strays from the original prompt. When an irrelevant response is given, we treat the model as skipping the question and assign it the worst possible sMAPE score (100%). To reduce the number of skipped questions, we introduce temperature backoff, which iteratively decreases the sampling temperature. Because the answer extractor can immediately identify when an output is incoherent, we can regenerate the response with a lower temperature to get the model to answer the task properly. Lower temperature increases coherence in VLMs [32], enhancing their chances of maintaining relevance to the prompt. Therefore, temperature backoff gives VLMs a better chance at achieving higher scores. Each time the answer extractor returns an empty answer, we reduce the temperature by (starting from ) until reaching a temperature of , at which point the example is skipped.
Models perform better with temperature backoff.
@@
We introduced temperature backoff to reduce model incoherence, and it performed fairly well. As shown in Tab. 6 (bottom), this method slightly improves performance across each model, resulting in an average error reduction of 5.78% in CAPTURe and 5.45% in CAPTURe. Temperature backoff essentially allows the model to reattempt the question if it fails to respond to the prompt. Similar to previous results, positive results from reattempts highlight VLMs’ weak reasoning abilities.
C.2 Only Occluded Object Counting
Models struggle at counting occluded objects.

### v2

@@ 4.2 Effect of Data Factors on VLM Performance
Here, we use the CAPTURe data (which can be controlled precisely and fully annotated ) to examine which features correlate with model performance. We test the effect of the following variables on final performance: (1) Increasing the number of occluded objects; (2) Varying the pattern. We also investigate whether models can classify patterns, and to what degree models can predict the number of occluded objects only (rather than the total).
Models perform worse when more dots are occluded. In Fig. 6 (right), we observe that error increases with respect to the number of occluded dots. However, Fig. 6 (left) also shows that performance is less affected by the total number of dots. This suggests that the task difficulty is more closely correlated with the difficulty of occlusion – i.e. the difficulty of the world modeling task – rather than the complexity of the pattern. Some models, such as GPT-4o, deviate from this trend, which has lower error on specific numbers. We further explore model bias in Sec. C.5.
Performance depends on pattern type. The controllability of CAPTURe allows us to measure the effect of pattern type on performance. In Fig. 7, we find that model performance differs across shapes with some regularity: objects arranged in a circle generally have lower sMAPE than other shapes, across all models.
@@
Models can identify patterns. To determine how much model errors can be attributed to a lack of pattern recognition ability, we formulate a separate task where models must recognize the pattern in the image on CAPTURe. Here, we frame the task as multiple-choice, asking the model to select from the pattern types available (rectangle, triangle, or circle). Table 4 illustrates that all perform substantially better than random at this task, with most models except InternVL2 achieving accuracy above 80% in the unoccluded setting. As expected, the patterns were easier to identify in unoccluded scenarios, with models suffering an average accuracy drop of 10.95% in the occluded setting. Notably, GPT-4o and Qwen2VL have a fairly small drop in performance, suggesting they can generally capture the pattern even in the presence of occlusion.
@@ 4.3 Analysis with Auxiliary Information
In Sec. 4.1, we see that models broadly struggle with amodal counting. Here, we seek to disentangle whether this problem results from a failure to reason, the absence of a world model, or both by giving VLMs two different types of auxiliary information: oracle information and predicted information. Oracle information is ground truth and is directly pulled from CAPTURe’s metadata, e.g., object locations. Predicted information generates new information from a completely separate model and gives it to the VLM. This information is not ground truth and is sourced from an external model, such as an image inpainting model, rather than the VLM. By giving the model auxiliary information in the form of reasoning and spatial clues, we can establish how much of each model’s error results from an inability to handle occlusion rather than an inability to recognize and count visible objects.
@@
Oracle setup. We test two oracles for CAPTURe’s occluded split based on its constituent subtasks: counting the visible objects and inferring/counting occluded objects. Both oracles provide the VLM with text-based coordinates of objects in the image, simplifying the visual task by assuming the VLM effectively has a perfect visual system that can recognize and localize objects in the image. The first oracle, the Visible Object Coordinate Oracle, gives the VLM the coordinates of all unoccluded objects (encoded as text, as seen in Fig. 8) and instructs the model to estimate the number of occluded objects, count the number of visible object coordinates, and add the two. In other words, the model is given oracle information about what objects are visible, thus also revealing key information about the pattern. The second oracle, the All Object Coordinate Oracle, instead gives the model the coordinates of all objects. Here, the model only needs to count the coordinates in the prompt, eliminating the need to reason on the visual input. Note that Molmo is excluded in these tests because it contains a prompt limit that would truncate the list of coordinates. An example of the oracle inputs can be seen in Fig. 8.
@@
Prediction setup. In this setting, we provide the VLM with an external world model representation predicted by another model. Specifically, we develop the Inpainting Pipeline to fill in the occluded region via a diffusion-based inpainting model and pass the inpainted image to the VLMs. For the inpainting model, we choose FLUX.1-Fill [dev], whose backbone FLUX.1 [dev] [21] is a top public model in the Text to Image Model Arena [7]. An example input to the VLM can be seen on the far-right of Fig. 8.
@@
Providing visible or all object coordinates improves performance substantially. The results in Tab. 2 indicate that models struggle on CAPTURe, which requires identifying a pattern and counting both visible and occluded objects. Moreover, models generally struggle with counting even in unoccluded settings. Both oracles simplify the counting task: All Object Coordinate Oracle reduces the task to simply counting coordinates with no reasoning involved, and Visible Object Coordinate Oracle similarly simplifies the task for visible objects, while still requiring inferring occluded objects. Additionally, under Visible Object Coordinate Oracle, recognizing the pattern shifts from a visual reasoning task to an augmented math problem. Instead of visually reasoning about where objects are located, the VLM considers what patterns the coordinates could make. Translating this task into a text problem results in an average increase of 15% with all objects coordinate oracle; the errors LLMs make here are due to an inability to count in the text prompt, as opposed to weaknesses in handling occlusion (since all object coordinates are given), and the strongest model, GPT-4o, achieves minimal error here. We also obtain an average increase of 8% with the visible objects coordinate oracle (shown in Tab. 5), possibly because it allows the more powerful LLM backbone (which is far larger than the vision model in all models tested) to complete the counting task. Taken together, these results suggest that there is much room for improvement in visual world modeling beyond text-based reasoning of VLMs.
@@
Providing diffusion-based inpainting improves performance marginally. Similar to the object coordinate oracles, the Inpainting Pipeline (rightmost columns in Fig. 8 and Tab. 5) eliminates the need for world modeling and provides VLMs with an approximation of the image behind the occluder. With the inpainted images, VLM error decreases by almost 2% for InternVL2 and 7% for Qwen2VL compared to the original occluded images. GPT-4o’s error increases on inpainted images by a small margin; we hypothesize that this may be because GPT-4o has one of the better world models (based on its superior performance), and thus does not improve further with the inpainted images. Moreover, every VLM still falls short of its unoccluded image performance, indicating that the diffusion model is not a perfect world model. Qualitatively, we find that the inpainting model sometimes fails to output the correct pattern.
@@ 5 Related Work
Spatial reasoning in visual question answering. Past work measures the spatial reasoning capabilities of VLMs in the form of visual question answering (VQA) [4, 16] benchmarks. SpartQA [26] asks VLMs to identify the spatial relation (e.g., above, behind, left of) between objects in synthetically created 2D images from NLVR [39]. More recent benchmarks test similar spatial relation understanding with real images [24, 36, 2]. While this past work asks models to provide a text description for a relation between two fully observed objects, CAPTURe measures the world modeling from a partially observed scene, thus requiring the handling of occlusion, pattern recognition, and counting. Together, these constitute a stricter test of spatial reasoning than typical VQA settings.
Amodal completion. Occlusions are common in natural scenes, and vision solutions for amodal completion have made significant progress in infilling occlusions [46, 38, 6]. The amodal completion task has evolved from simply completing a shape to filling in appearance (e.g., texture, color, etc.) to finally dealing with fine-grained order perception (multiple stacked occluded objects) [5]. Specifically in Qiu and Di [34], VLMs classify the hidden objects and extract fine details from occluded items. CAPTURe, however, presents a unique category of patterned amodal counting which requires inferring fully occluded objects based on a pattern rather than inferring occluded object wholes based on object parts. In other words, previous work has only attempted tasks that require amodal completion for one object at a time [46, 31, 38], whereas CAPTURe handles multiple objects. Multi-object amodal completion is crucial because in cluttered scenes, entire groups of objects are often occluded. Moreover, the output space of CAPTURe is language (rather than filling pixels).
Counting with vision-and-language models. Within the task of counting, the most similar application to CAPTURe is dense counting, where the objects to be counted occlude each other. There are many practical applications of such a task, like counting cells on a crowded slide [8], determining crop yields from densely-packed fields [43], or crowd counting [44, 48, 14]. Liang et al. [23] improved crowd counting with an augmented CLIP [35], i.e. also using VLMs for counting. Additionally, Jenkins et al. [18] introduced an amodal counting benchmark, presenting an occluded 3D counting task where models must count objects on retail shelves. However, our work differs in many ways, as Jenkins et al. [18] only counts retail shelves and uses LiDAR input. More broadly, dense counting focuses on overlapping objects rather than on counting objects arranged into patterns, which is the focus of CAPTURe.
@@ 6 Conclusion
We introduced CAPTURe, a novel benchmark for amodal counting that measures spatial reasoning capabilities under occlusion. CAPTURe is designed to assess VLMs’ ability to form a robust world model and use that model for visual reasoning skills under occlusion. By testing counting, we cast the problem as a measurable task with an objective correct answer that also has real-world utility as VLMs become more broadly adopted. Our results suggest that VLMs struggle to combine reasoning, counting, and world modeling with low performance on occluded and unoccluded images. Our analysis indicates that models improve with oracle information about visible objects (simplifying the reasoning/counting tasks) and predicted information about the occluded objects (also simplifying world modeling), pointing to directions of model improvement.
@@ A.1 Metric Details
We use symmetric mean percent error (sMAPE) as the primary metric for our benchmarks due to its resistance to bias for under/over predictions and small/large ground truths [25]. The standard metric for a counting benchmark is mean average error (MAE). MAE is popular, but heavily penalizes predictions that deviate by a small margin from big ground truths, highlighting the necessity for a metric that gives equal weighting to all questions. Mean average percent error (MAPE) initially seems appealing but is disproportionally inflated for small ground truths and is biased towards overpredictions. Mean square error (MSE) and root mean square error (RMSE) are also commonly used but are very sensitive to outliers because they square the error. Intuitively, performing well on almost all questions and poorly on a small subset should score better than consistently being wrong. Among commonly-used metrics, sMAPE is the only metric that evaluates performance in relation to the distribution of ground truth elements [11]. There are two common definitions [15] for sMAPE, but we use the one that scales to 100%. sMAPE is given by:
@@ A.2 Output Tokens
To maximize the VLM’s chance at success, we allocate a high number of output tokens to generate a rationale and output. This varies per model. We give 4000 tokens to InternVL2, 2000 tokens to Molmo, and 8192 tokens to Qwen2VL, following their max output lengths. For GPT-4o, we use the default of 4096 tokens.
@@ Appendix B CAPTURe Dataset Creation Details
The following expands upon Sec. 2.2. While FSC-147, a diverse counting dataset with manual annotations, is a strong starting point, it cannot immediately be adapted to our task. To make the task of amodal counting solvable, our dataset requires images with patterns in them. A person (or model) can infer how the pattern would continue and thus accurately predict the total number. For questions to be answerable, the dataset’s images must be filtered down to represent patterns a model or person could recognize.
Our filtering process follows two stages. First, we prompt GPT-4o to determine whether the objects were arranged in a pattern. Second, if the model responded with “no”, the images were immediately discarded. If the model output was “yes”, the log probability of the token is stored. Empirically, we found that higher log probability values (i.e. higher confidence scores) corresponded to more well-defined patterns in the image. Thus, we use the log probabilities for filtering.
Specifically, let be the log probability of the “yes” token and denote the threshold for determining how well-defined a pattern is.222We set based on manual evaluation, finding it resulted in fewer false positives. To filter the images based on pattern rigidity, we apply the following condition: . This inequality yields 991 images from the original dataset (16.12%). Next, we manually filter each of the selected images to ensure that they indeed contain patterns and feature a countable number of objects, excluding 34 images. Afterward, we manually place a “fair” occluding box in each image, i.e. a box that leaves sufficient portions of the pattern visible, such that the pattern can still be inferred from the unoccluded portions of the image. Occluding boxes were also chosen with varying positions and sizes in the image.
@@ Appendix C Additional Analysis
Here we provide additional experiments that attempt to either increase model performance on CAPTURe or dissect the reasons behind poor model performance. Chain-of-Thought inhibits model performance, while temperature backoff slightly improves performance. Additionally, we find that models struggle at counting just occluded objects, are overconfident in occluded settings, and are biased to predict specific numbers.
C.1 Chain-of-Thought reduces model performance
Method | CAPTURe | CAPTURe |
GPT-4o | 14.75 | 9.71 |
GPT-4o w/ CoT | 14.94 | 7.73 |
Qwen2 | 29.33 | 11.74 |
Qwen2 w/ CoT | 31.57 | 37.81 |
During development, we experimented with several common strategies including CoT. In Tab. 6, we find that CoT reduces model performance except in the occluded synthetic scenario, most likely because the included examples are very similar to the test prompt.
C.2 Temperature backoff slightly improves model performance
Model | Error (%) () | |||||||
@@ C.2 Temperature backoff slightly improves model performance
Unoccluded | Occluded | Unoccluded | Occluded | |||||
@@
GPT-4o | 13.34 | 12.57 () | 14.75 | 14.39 () | 5.90 | 5.93 () | 9.71 | 9.23 () |
InternVL2 | 26.17 | 27.09 () | 32.90 | 32.37 () | 16.44 | 15.59 () | 17.57 | 16.24 () |
Molmo | 25.90 | 21.23 () | 32.49 | 28.17 () | 8.40 | 2.88 () | 17.73 | 15.85 () |
Qwen2VL | 18.96 | 19.40 () | 29.33 | 28.47 () | 6.63 | 6.66 () | 11.74 | 11.51 () |
Avg. of 4 VLMs | 21.09 | 20.07 () | 27.37 | 25.85 () | 9.34 | 7.76 () | 14.19 | 13.21 () |
To improve VLM performance on CAPTURe, we address a trend we established during early testing. Most of the time, the VLM fails by reaching an incorrect answer. Sometimes, however, our benchmark can cause VLMs to produce a long and irrelevant response that strays from the original prompt, leading to the worst possible sMAPE score (100%).
@@
To reduce the number of skipped questions, we experiment with temperature backoff, which iteratively decreases the sampling temperature. Because the answer extractor can immediately identify an incoherent output, we can regenerate the response with a lower temperature to get the model to answer the task properly. Consistent with our findings, Peeperkorn et al. [32] also finds that lower temperatures increase coherence in VLMs, thereby enhancing their chances of maintaining relevance to the prompt. Therefore, temperature backoff gives VLMs a better chance of achieving higher scores. Each time the answer extractor returns an empty answer because the VLMs produced an incoherent answer, we reduce the temperature by (starting from ) until it reaches , at which point the example is skipped.
@@
Models perform slightly better with temperature backoff. We introduced temperature backoff to reduce model incoherence, and it performed fairly well. As shown in Tab. 7 (bottom), this method slightly improves performance across each model, resulting in an average error reduction of 5.78% in CAPTURe and 5.45% in CAPTURe. Temperature backoff essentially allows the model to reattempt the question if it fails to respond to the prompt. Similar to previous results, positive results from reattempts highlight VLMs’ weak reasoning abilities.
C.3 Models struggle at counting just occluded objects